Find two numbers whose sum is 39 and such that one is 2 times as large as the other. (We will call the two numbers x and y.)
Answer by ptaylor(2198)


Let x=one number
Then 39-x=the other number
Now we are told that one number (lets say x) is 2 times as large as the other, so our equatin to solve is:
x=2(39-x) get rid of parens
x=78-2x add 2x to each side
x+2x=78-2x+2x collect like terms
3x=78 divide each side by 3
x=26---------------------------one number
39-x=39-26=13---------------the other number

OK now we will call the two numbers x and y
x+y=39-------------------eq1
and
y=2x----------------eq2
substitute y=2x into eq1
x+2x=39
3x=39
x=13 one number
y=2x=2*13=26 the other number

Does this help???-------------ptaylor